### v1

# Scaling and Predictability in Surface Quasi-Geostrophic Turbulence

Abstract
Turbulent flows are strongly chaotic and unpredictable, with a Lyapunov exponent that increases with the Reynolds number. Here, we study the chaoticity of the Surface Quasi-geostrophic system, a two-dimensional model for geophysical flows that displays a direct cascade similar to that of three-dimensional turbulence. Using high-resolution direct numerical simulations, we investigate the dependence of the Lyapunov exponent on the Reynolds number and find an anomalous scaling exponent larger than the one predicted by dimensional arguments. We also study the finite-time fluctuation of the Lyapunov exponent by computing the Cramér function associated with its probability distribution. We find that the Cramér function attains a self-similar form at large Re.
keywords:
Surface Quasi-geostrophic, Turbulence, Predictability, Direct Numerical Simulations.
1 Introduction
Turbulence is a complex and chaotic phenomenon characterized by a large number of interacting degrees of freedom organized hierarchically across multiple scales of motion (Frisch, 1995). Determining whether a turbulent flow remains predictable or, at each scale, retains any degree of predictability has been a longstanding challenge, tracing back to the pioneering works by Lorenz, Ruelle, Leith, and Kraichnan (Lorenz, 1969; Leith, 1971; Leith & Kraichnan, 1972; Ruelle, 1979; Deissler, 1986). The nonlinear amplification of small-scale perturbations led to the formulation of the famous “butterfly effect” (Lorenz, 1963). When extended to multiscale systems such as turbulence, these ideas give rise to the concept of the cascade of errors, in which small-scale perturbations progressively amplify and propagate to larger scales, gradually spoiling predictability at larger and larger scales (Boffetta et al., 1997; Rotunno & Snyder, 2008; Palmer, 2024; Boffetta & Musacchio, 2017).
A key mathematical tool for studying predictability is the Lyapunov exponent and its finite-time version (FTLE). Ruelle predicted that the Lyapunov exponent in turbulence is proportional to the inverse of the smallest time, the Kolmogorov time, and therefore, it increases with the flow’s Reynolds number (Re) (Ruelle, 1979). Nonetheless, a turbulent flow at large Re remains predictable at large scales since the error grows with the characteristic turnover time of that scale, which is independent of the Reynolds number. This property is familiar in oceanography and atmospheric flows where the smallest time scale can be very small (fraction of seconds) (Garratt, 1994) while the weather remains predictable for days.
In this work, we consider the surface quasi-geostrophic (SQG) equation, a model that describes the flow governed by the conservation of buoyancy at the surface of a rotating, stratified fluid (Blumen, 1978; Pierrehumbert et al., 1994). Beyond its geophysical relevance, the SQG model has gained attention in the fluid dynamics community due to its striking similarities to three-dimensional (3D) Navier-Stokes (NS) turbulence while keeping some properties of two-dimensional (2D) flows. In particular, SQG has two inviscid quadratic invariants, similarly to 2D turbulence (Celani et al., 2004; Lapeyre, 2017; Valade et al., 2024) with one of the two, the surface potential energy, which displays, in the presence of forcing and dissipation, a direct cascade à la Kolmogorov towards the small scales similar to 3D turbulence (Valadão et al., 2024b). Despite this similarity, previous numerical studies (Pierrehumbert et al., 1994; Ohkitani & Yamada, 1997; Sukhatme & Pierrehumbert, 2002) reported that the scaling exponent of the spectrum of the surface potential energy deviates from the Kolmogorov value $-5/3$ predicted by dimensional arguments.
Based on very high-resolution direct numerical simulations of the SQG model, we study the statistical properties of the direct cascade. At large Reynolds number, we observe the recovery of the Kolmogorov scaling in the surface potential energy spectrum. Further, we measure the finite-time distribution of the Lyapunov exponent as a function of Re and we find an anomalous scaling law in which the Lyapunov exponent grows faster than what is predicted on dimensional grounds, similar to what was observed in 3D NS turbulence (Boffetta & Musacchio, 2017). Despite this anomaly, we find that the distribution of the finite-time Lyapunov exponents follow an almost universal function, independent of the Reynolds number of the flow.
The remainder of this paper is as follows. In section 2 we review the basic definitions of the SQG equation and discuss its statistical properties in the turbulent regime. Section 3 discusses the main results obtained in this work through the use of extensive numerical simulations covering a large range of Reynolds numbers. We split the results into two subsections: Section 3.1 explores the Reynolds dependence on the dimensional scaling properties of SQG, principally the scaling exponent of the surface potential energy spectrum; Section 3.2 addresses the Eulerian predictability and the statistics of finite-time Lyapunov exponent (FTLE) as functions of Reynolds number. In Section 4, we summarize our results, pointing out directions for future research.
2 The Surface Quasi-geostrophic model
The SQG equation describes the evolution of the surface buoyancy field $\theta({\bm{x}},t)$, governed by the following equation (Pierrehumbert et al., 1994):
$\partial_{t}\theta+{\bf v}\cdot{\bm{\nabla}}\theta=\nu\nabla^{2}\theta-\mu\nabla^{-2}\theta+f\;,$ (1)
where $f(\bm{x},t)$ represents a forcing, $\nu$ is a diffusion coefficient, and $\mu$ accounts for large-scale friction. We remark that most oceanic and atmospheric applications of SQG disregard dissipation terms as the model is designed to capture large-scale dynamics (Juckes, 1994; Lapeyre & Klein, 2006; Vallis, 2017; Siegelman et al., 2022). On the contrary, when studied numerically as a turbulent system, the model requires dissipative terms to avoid the small and large scale accumulation of the energy injected by the forcing term and to reach a statistically stationary state.
The incompressibility condition $\bm{\nabla}\cdot\bf v=0$ can be enforced through the definition of a stream function $\psi(\bm{x},t)$, such that ${\bf v}({\bm{x}},t)=(-\nabla_{y}\psi,\nabla_{x}\psi)$. Finally, the relation between the buoyancy field and the stream function is given by $\psi=\nabla^{-1}\theta$ or, in Fourier space, $\hat{\psi}=\hat{\theta}/k$ (where $k\equiv|\bm{k}|$) and consequently the velocity field can be written in terms of buoyancy as
$\hat{\bf v}({\bm{k}})=\left(-\frac{ik_{y}}{k},\frac{ik_{x}}{k}\right)\hat{\theta}({\bm{k}})$ (2)
from which one observes that the buoyancy field is dimensionally a velocity.
In the absence of forcing and dissipation ($f=0$, $\nu=0$, $\mu=0$), the SQG equation (1) conserves two quadratic quantities, the vertically integrated energy (VIE),
$V=\frac{1}{2}\langle\psi\theta\rangle\;,$ (3)
and the surface potential energy (SPE),
$E=\frac{1}{2}\langle\theta^{2}\rangle\;,$ (4)
where brackets stand for spatial average.
If one assumes that the forcing $f$ is active on a characteristic scale $\ell_{f}$, a turbulent state can develop with a double cascade phenomenology (Blumen, 1978; Pierrehumbert et al., 1994). Within this scenario, SPE is primarily transferred from the forcing scale to smaller ones ($\ell<\ell_{f}$), producing the direct cascade which is eventually dissipated by viscosity at the diffusive scale $\ell_{\nu}$. Meanwhile, VIE undergoes an inverse cascade, transferring energy to scales larger than the forcing ($\ell>\ell_{f}$) until it is dissipated at the friction scale $\ell_{\mu}$. In the statistically stationary state, the SPE and VIE balances are given by
$\varepsilon_{I}=\varepsilon_{\nu}+\varepsilon_{\mu}\;,$ (5)
$\eta_{I}=\eta_{\nu}+\eta_{\mu}\;,$ (6)
where $\varepsilon_{I}=\left\langle\theta f\right\rangle$ and $\eta_{I}=\left\langle\psi f\right\rangle$ are the input rates, $\varepsilon_{\nu}=\nu\left\langle|{\bm{\nabla}}\theta|^{2}\right\rangle$ and $\eta_{\nu}=\nu\left\langle\bm{\nabla}\psi\cdot\bm{\nabla}\theta\right\rangle$ the small-scale dissipation rates while $\varepsilon_{\mu}=\mu\left\langle\theta\nabla^{-2}\theta\right\rangle$ and $\eta_{\mu}=\mu\left\langle\psi\nabla^{-2}\theta\right\rangle$ are the large-scale dissipation rates of SPE and VIE, respectively.
Under the assumptions of statistical homogeneity and isotropy, Blumen (Blumen, 1978) first predicted the power-law behavior of the energy spectrum $E(k)\equiv\langle|\hat{\theta}({\bm{k}})|^{2}\rangle/2$ for sufficiently large scale separations $\ell_{\nu}\ll\ell_{f}\ll\ell_{\mu}$. In this case, the spectral energy densities follow
$E(k)\simeq\eta_{\mu}^{2/3}k^{-1},\quad 1/\ell_{\mu}\ll k\ll 1/\ell_{f}$ (7)
$E(k)\simeq\varepsilon_{\nu}^{2/3}k^{-5/3},\quad 1/\ell_{f}\ll k\ll 1/\ell_{\nu}$ (8)
and the diffusive and friction scales are determined on dimensional grounds as
$\ell_{\nu}\equiv\left(\frac{\nu^{3}}{\varepsilon_{\nu}}\right)^{1/4},\quad\ell_{\mu}\equiv\left(\frac{\eta_{\mu}}{\mu^{3}}\right)^{1/9}.$ (9)
The ratio between the dissipative and forcing scales, i.e., the extension of the inertial range, defines the Reynolds numbers associated with the flow. For the direct cascade of SPE, in analogy to 3D NS turbulence, we define the Reynolds number as
$\text{Re}\equiv\frac{\varepsilon_{I}^{1/3}\ell_{f}^{4/3}}{\nu}\simeq\left(\frac{\ell_{f}}{\ell_{\nu}}\right)^{4/3}$ (10)
Note that in (10) Re is based on $\varepsilon_{I}$ and it is therefore defined a priori. Alternatively, we could use the VIE dissipation $\varepsilon_{\nu}$ resulting in a slightly smaller value of Re, as discussed below. We also remark that the scaling laws in (7-8) can also be obtained from the analogous of the exact four-fifths law of turbulence and by assuming self-similarity of the statistics (Frisch, 1995; Valadão et al., 2024b; Valade et al., 2025).
3 Numerical simulations and results
We explore the statistical properties and the Eulerian predictability of the direct cascade in SQG at different Reynolds numbers by numerically integrating (1) at high resolution with a pseudo-spectral, GPU accelerated code. Simulations are performed in a square domain of size $L_{x}=L_{y}=2\pi$ with periodic boundary conditions, using a regular grid with resolution $N\times N$ ranging from $N=1024$ to $N=16384$. Simulations cover more than two decades in the diffusion coefficient, corresponding to a Reynolds number (10) which varies from $\text{Re}=600$ to $\text{Re}=158800$, while the large scale dissipation coefficient $\mu$ is fixed. The system is driven by a constant-amplitude forcing with random phases, active within a narrow circular shell in wavevector space centered on $k_{f}$ and with a small width $\Delta k$. This forcing provides constant SPE and VIE injection rates $\varepsilon_{I}$ and $\eta_{I}$ respectively with $\varepsilon_{I}\approx\eta_{I}k_{f}$ since $\Delta k\ll k_{f}$. Specific details on the GPU code performances can be found in (Valadão et al., 2024a).
The most relevant parameters on the simulations are listed in Table 1. All the simulations are performed in statistically stationary states, including the subset of simulations for computing the Lyapunov exponent. We also performed a careful study on the sensitivity of the following results to the maximum resolved wavenumber $k_{\max}\ell_{\nu}$ by increasing resolution at fixed Re. We found independence of the results on the resolution for $k_{\max}\ell_{\nu}\gtrsim 1.5$.
Run $N$ Re $k_{\max}\ell_{\nu}$ $\tau_{\nu}$ $T_{tot}/\tau_{f}$ Run $N$ Re $k_{\max}\ell_{\nu}$ $\tau_{\nu}$ $T_{tot}/\tau_{f}$
$A_{1}$ 1024 $642$ $5.5$ $0.0260$- $C_{2}$ 4096 $10600$ $2.9$ $0.0075$ $1130$
$A_{2}$ 1024 $794$ $4.7$ $0.0240$- $C_{3}$ 4096 $15900$ $2.1$ $0.0062$ $1130$
$A_{3}$ 1024 $1060$ $3.8$ $0.0210$- $C_{4}$ 4096 $21200$ $1.7$ $0.0054$ $1130$
$A_{4}$ 1024 $1590$ $2.9$ $0.0180$- $C_{5}$ 4096 $25400$ $1.5$ $0.0049$ $1130$
$B_{1}$ 2048 $2120$ $4.7$ $0.0160$- $D_{1}$ 8192 $31800$ $2.5$ $0.0044$ $9\times 287$
$B_{2}$ 2048 $3180$ $3.5$ $0.0130$ $2270$ $D_{2}$ 8192 $42400$ $2.1$ $0.0039$ $9\times 287$
$B_{3}$ 2048 $6350$ $2.1$ $0.0099$- $D_{3}$ 8192 $63500$ $1.5$ $0.0031$ $9\times 287$
$B_{4}$ 2048 $7940$ $1.8$ $0.0086$ $2270$ $E_{1}$ 16384 $90800$ $2.3$ $0.0026$-
$C_{1}$ 4096 $3970$ $5.9$ $0.0120$- $E_{2}$ 16384 $159000$ $1.6$ $0.0020$-
Table 1: Relevant parameters of the simulation: Reynolds number, diffusive scale $\ell_{\nu}=\nu^{3/4}\varepsilon_{\nu}^{-1/4}$, diffusive time $\tau_{\nu}=\sqrt{\nu/\varepsilon_{\nu}}$, total length of the Lyapunov simulations $T_{\text{tot}}$. Common parameters for all simulations: forcing wavenumber $k_{f}=3.5$ and width $\Delta k_{f}=0.5$, surface potential energy input $\varepsilon_{I}=24$, friction coefficient $\mu=1.0$, characteristic time at the forcing scale $\tau_{f}=\varepsilon_{I}^{-1/3}\ell_{f}^{2/3}=0.51$, maximum resolved wavenumber $k_{\max}=N/3$ (2/3 dealiasing rule).
3.1 Statistics of the direct cascade
For large separations between the forcing and the dissipative scales, one expects that almost all the SPE is transferred and dissipated at small scales, while VIE is dissipated at large scales. This is the essence of the argument developed by Fjortoft for 2D turbulence (Fjørtoft, 1953) and verified by numerical simulations of 2D NS double cascade (Boffetta & Musacchio, 2010). In the present case, using (5) and (6) and the scaling relation $\eta_{\ell}\simeq\ell\varepsilon_{\ell}$ this argument gives
$\frac{\eta_{\nu}}{\eta_{\mu}}=\left(\frac{\ell_{\mu}-\ell_{f}}{\ell_{f}-\ell_{\nu}}\right)\frac{\ell_{\nu}}{\ell_{f}}\simeq\text{Re}^{-3/4}$ (11)
where we have used (10) to express $\ell_{\nu}/\ell_{f}$ as a function of Re.
Figure 1: Main figure: SPE relative dissipation $\varepsilon_{\nu}/\varepsilon_{I}$ (full circles) and VIE relative dissipation $\eta_{\nu}/\eta_{I}$ (open circles), both as functions of Re. The inset shows $\eta_{\nu}/\eta_{I}$ as functions of Re on a log-log scale.
Figure 1 shows the fraction of small-scale dissipation of the inviscid invariants as a function of the Reynolds number of the flow in statistically stationary conditions. Indeed, we find that the small scale relative dissipation of VIE vanishes in the limit of large Re following the prediction (11). On the contrary, since we do not resolve the inverse cascade and $\ell_{\mu}\sim\ell_{f}$, there remains a constant large-scale SPE dissipation $\varepsilon_{\mu}$ for large Re. Therefore, the direct cascade transfers only a fraction of the total injected energy equivalent to $\varepsilon_{\nu}\approx 0.45\varepsilon_{I}$ for $\text{Re}\gtrsim 10^{4}$.
Figure 2: Normalized SPE direct cascade fluxes $\Pi(k)$ of all runs on table 1. The dashed line represents $\varepsilon_{\nu}/\varepsilon_{I}=0.45$.
Stationary fluxes of SPE in Fourier space are presented in Fig. 2 for different Reynolds numbers. At moderate $\text{Re}\lesssim 3000$ the fluxes for $k>k_{f}$ decay quickly as a consequence of the viscous dissipation. For large Re, however, a plateau of constant flux emerges at a level corresponding to the viscous dissipation rate $\varepsilon_{\nu}$. We emphasize that SQG turbulence exhibits large fluctuations in the flux of the direct cascade, as studied in details in (Valadão et al., 2024b). These fluctuations arise from the interplay between the accumulation of energy in large-scale structures and intense dissipative events triggered by the formation of filamentary shocks that transfer energy from large to small scales over short time intervals. Thus, very long integrations are necessary to observe the convergence to the constant flux plateau of Fig. 2.
Figure 3: Main plot: Time-averaged spectra $E(k)$ for all simulations. Color coding follows the same as Fig. 2. Inset: $C_{K}$ as functions of the Reynolds number.
The time-averaged spectra $E(k)$ of SPE are shown in Fig. 3 for all the runs in Table 1. All spectra exhibit power-law behavior, $E(k)\propto k^{-\beta}$ in an intermediate range of scales, which becomes wider as Re increases. We observe that at moderate Re, when a scaling range is already clearly observable, the scaling exponent $\beta$ deviates significantly from the dimensional prediction $5/3$, a feature already reported by previous investigations at comparable Reynolds numbers (Pierrehumbert et al., 1994; Ohkitani & Yamada, 1997; Sukhatme & Pierrehumbert, 2002). Nonetheless, we find that when Re is sufficiently large, $\text{Re}\gtrsim 10^{4}$, the scaling of the dimensional prediction (8) is closely recovered.
To quantify this important result, we measured the correction $\xi$ to the dimensional scaling exponent by fitting the intermediate range of the spectra with
$E_{K}(k)=C_{K}\varepsilon_{\nu}^{2/3}k^{-5/3}\left(\frac{k}{k_{f}}\right)^{-\xi}$ (12)
where $\xi$ and $C_{K}$ are the fitting parameters. In order to estimate the robustness of the fit, we adopted the following procedure: For each run, we fit the data with (12) in a range of wavenumbers $k\in[k_{0},k_{1}]$ with varying $k_{0}\in[3k_{f},5k_{f}]$ and $k_{1}\in[8k_{f},10k_{f}]$. This produces a set of parameters $\xi$ and $C_{K}$ for each run, from which we compute the mean using twice the standard deviation as an estimation of the error.
Figure 4: Main plot: scaling exponent correction $\xi$ to the SPE spectrum as a functions of Re. Inset: $E(k)$ compensated with $E_{K}(k)$ given by (12). The runs follow the same color coding as in Fig. 2.
In Fig. 4, we plot the dependence of $\xi$ on Re, together with the spectra of Fig. 3 compensated with the expression (12). It is evident that, while for $\text{Re}\lesssim 10^{4}$, the exponent correction $\xi$ depends on Re (approximately as $\text{Re}^{-3/4}$), for larger values of Re, the correction decreases much faster and becomes smaller than $5\%$ for $\text{Re}>2\times 10^{4}$. In this limit, we also observe the convergence of the dimensionless constant to $C_{K}=5.05\pm 0.11$ (see Fig. 3).
We remark that this behavior, which suggests the existence of a minimum Reynolds number for the recovery of dimensional scaling is very different to what was observed in the direct cascade of 3D NS turbulence, where Kolmogorov scaling is observed as soon as the spectrum displays a power-law behavior.
3.2 Predictability of the direct cascade
For the study of chaos and predictability of the direct cascade of SPE, we are interested in computing how two solutions $\theta(\bm{x},t)$ and $\theta^{\prime}(\bm{x},t)$ separate in time on average. In this paper, we consider infinitesimally close solutions so that the average separation rate is given by the maximal Lyapunov exponent of the flow.
Starting from a solution $\theta(\bm{x},t)$ of (1) in a statistically stationary state, we generate a perturbed solution as $\theta^{\prime}(\bm{x},t)=\theta(\bm{x},t)+2\sqrt{\Delta}W({\bm{x}})$ where $W({\bm{x}})$ is a Gaussian random noise with zero mean and unit variance while $\Delta$ is a small parameter. The SPE error $E_{\Delta}$ is defined, for any time, as
$E_{\Delta}(t)={1\over 2}\langle\delta\theta({\bm{x}},t)^{2}\rangle$ (13)
where the difference field is $\delta\theta=(\theta^{\prime}-\theta)/\sqrt{2}$ and the normalization coefficient $1/\sqrt{2}$ ensures that $E_{\Delta}=E$ for two completely uncorrelated fields. At initial time, by definition, we have $E_{\Delta}(t)=\Delta$. We measure the finite-time Lyapunov exponent by computing the growth rate of the error
$\gamma_{\tau}(t)={1\over 2\tau}\ln\left({E_{\Delta}(t+\tau)\over\Delta}\right)$ (14)
and then by rescaling the perturbed field to the initial SPE error
$\theta^{\prime}\leftarrow\theta-\sqrt{\frac{\Delta}{E_{\Delta}}}\left(\theta-\theta^{\prime}\right)\,.$ (15)
By repeating the steps (14) and (15) over many time intervals of the same length $\tau$, we obtain a distribution of FTLE along the trajectory. The rescaling procedure (15) ensures the permanence of the perturbation in the exponential growth regime when $\Delta$ and $\tau$ are sufficiently small.
From the definition (14) one can compute the FTLE for any time multiple of $\tau$, $T=n\tau$, simply by averaging
$\gamma_{T}(t)={1\over n}\sum_{k=1}^{n}\gamma_{\tau}(t+k\tau)$ (16)
and the Lyapunov exponent is given by the average of FTLE over a very long trajectory (and become independent of the initial condition)
$\lambda=\lim_{T\to\infty}\gamma_{T}(t)\,.$ (17)
In general, the distribution of FTLE around the Lyapunov exponent, for sufficiently large $T$, follows the large deviation principle (Vulpiani et al., 2009) which states that
$\rho(\gamma_{T})={1\over N_{T}}e^{-TC(\gamma_{T})}$ (18)
where $N_{T}$ is a normalizing factor and $C(\gamma_{T})$ is the Cramér function, independent on $T$ which, in general, vanishes at $\gamma_{T}=\lambda$ and is positive for $\gamma_{T}\neq\lambda$. For not too large fluctuations, the Cramér function can be approximated by a quadratic form
$C(\gamma_{T})\approx\frac{(\gamma_{T}-\lambda)^{2}}{2\Omega}\ .$ (19)
where $\Omega$, proportional to the variance of the distribution $\rho(\gamma_{T})$, is obtained from
$\Omega=\lim_{T\to\infty}T\langle(\gamma_{T}-\lambda)^{2}\rangle$ (20)
and it is expected to be $T$-independent in the limit of large $T$.
Figure 5: Fields $\theta(\bm{x})$ and $\delta\theta(\bm{x})$ of Run $D_{3}$ are shown in the left and right panels, respectively. The color code of $\theta(\bm{x})$ shows white/black as the most negative/positive value. The right panel’s color code is based on a log scale $(\propto\log_{10}(|\delta\theta(\bm{x})|))$ to facilitate visualization.
We computed the FTLE for simulations at different Reynolds numbers corresponding to runs $B_{2}$, $B_{4}$, $C_{2}$, $C_{3}$, $C_{4}$, $C_{5}$, $D_{1}$, $D_{2}$, and $D_{3}$ of Table 1. For all runs, we excluded the initial FTLEs in Eq. (14) from the statistics, allowing the perturbation to align with the most unstable direction of the system. For the three cases at the highest Reynolds numbers, we compensated for the increased computational cost by averaging over nine independent, shorter simulations run in parallel, each of which with different realizations of the forcing $f(\bm{x})$ and initial perturbation noise $W(\bm{x})$.
Figure 5 shows a representative realization of the field $\theta(\bm{x})$ along with its corresponding perturbation field $\delta\theta(\bm{x})$. The error seems to accumulate predominantly in filamentary zones between coherent structures. These regions are dominated by small-scale structures formed by energy transfer from larger scales (Pierrehumbert et al., 1994). Since such structures appear intermittently in time, the convergence of the FTLE statistics requires very long simulations, as discussed below.
Figure 6 presents the FTLE for the different runs as a function of the average time $T$. We see that in all the cases, the average FTLE converges, after a long transient and for $T\gtrsim 200\tau_{f}$, to the asymptotic value, which represents the Lyapunov exponent of the flow.
Figure 6: Convergence of the FTLE as a function of the average time. Light to dark colors represents runs at increasing Reynolds numbers $B_{2}$, $B_{4}$, $C_{2}$, $C_{3}$, $C_{4}$, $C_{5}$, $D_{1}$, $D_{2}$ and $D_{3}$.
From Fig. 6, it is evident that the Lyapunov exponent increases with the Reynolds number. By adapting the dimensional arguments developed for 3D NS turbulence (Ruelle, 1979), we expect that $\lambda$ is proportional to the inverse of the smallest dynamical time, i.e., the Kolmogorov time $\tau_{\nu}\equiv(\nu/\varepsilon_{\nu})^{1/2}$
$\lambda\simeq{1\over\tau_{\nu}}\simeq{1\over\tau_{f}}\text{Re}^{1/2}\ .$ (21)
Figure 7: Reynolds scaling of the mean FTLE $\left\langle\gamma_{T}\right\rangle=\lambda$. On the left panel, nondimensionalization is made with $\tau_{f}$, while the right panel uses $\tau_{\nu}$.
In Fig. 7, we plot the Lyapunov exponents of our simulations as functions of Re. We find that $\lambda$ grows with Re faster than what predicted by (21) and the best fit gives $\lambda\tau_{f}\simeq\text{Re}^{0.7}$ or, equivalently, $\lambda\tau_{\nu}\simeq\text{Re}^{0.2}$. In Fig. 7 we also plot the scaled variance $\Omega$ as a function of Re, which displays a scaling law compatible with that of the Lyapunov exponent $\Omega\tau_{f}\simeq\text{Re}^{0.7}$. This indicates that, in the range of Re investigated here, the ratio $\Omega/\lambda$ is approximately constant ($0.23\pm 0.02$) and that the central part of the Cramér function has a self-similar evolution with Re.
A qualitatively similar behavior has been observed for the Lyapunov exponent of 3D turbulence (Boffetta & Musacchio, 2017; Berera & Ho, 2018; Ge et al., 2023) but with a correction to the dimensional scaling (21) smaller than in the present case. We remark that the origin of such a correction, in both cases, cannot be attributed to intermittent fluctuations in the cascade, which are present also in the SQG cascade of SPE (Valade et al., 2025), since intermittency correction would require an exponent which is smaller than the dimensional one predicted by (21) (Aurell et al., 1996).
Figure 8: Cramér function for the runs at different Re normalized with their corresponding Lyapunov exponent $\lambda$. The dashed line represents the quadratic form $(x-1)^{2}/(2\Omega/\lambda)$ with $\Omega/\lambda=0.23$.
In figure 8, we show the Cramér function $C(\gamma_{T})$ for the different runs obtained from Eq. (18) at sufficiently large values of $T\lambda$. Since $\lambda$ and $\Omega$ scale with Re with the same exponent, we rescale both $\gamma_{T}$ and $C(\gamma_{T})$ by $\lambda$ which, according to the quadratic approximation (19), predicts the collapse to the function $(x-1)^{2}/(2\Omega/\lambda)$. From Fig. 8, we see a remarkable collapse of the Cramér functions at different Re even if the fluctuation for large positive $\gamma_{T}$ is far from the Gaussian prediction (18).
4 Conclusions
In this study, we investigated the statistical properties and predictability of turbulence in the Surface Quasi-Geostrophic (SQG) model using high-resolution direct numerical simulations across a wide range of Reynolds numbers. Our analysis focused on two central aspects: the scaling behavior of the energy spectrum in the direct cascade of surface potential energy (SPE) and the chaotic dynamics characterized by finite-time Lyapunov exponents (FTLEs).
Our results indicate that, for $\mathrm{Re}\gtrsim 2\times 10^{4}$, the energy spectrum approaches the Kolmogorov-like scaling $E(k)\propto k^{-5/3}$. This observation, together with the convergence in Reynolds of the prefactor $C_{K}$ suggests that SQG turbulence exhibits a well-defined inertial range—similar to that of 3D Navier-Stokes turbulence, but it does so only at very large Reynolds numbers. This regime was not observed by earlier studies at moderate Reynolds numbers.
In terms of predictability, we showed that the Lyapunov exponent scales anomalously with the Reynolds number as $\lambda\propto\text{Re}^{0.7}$, exceeding the dimensional prediction $\lambda\propto\text{Re}^{1/2}$. This anomalous scaling is reminiscent of similar behavior observed in 3D turbulence, where $\lambda\propto\text{Re}^{0.64}$, suggesting that in both systems, the deviation from dimensionality originates from a shared dynamical property. Beyond the average growth rate of infinitesimal perturbations, we investigated the statistical properties of FTLEs. Notably, both the variance and the shape of the associated Cramér function exhibit self-similar behavior across Reynolds numbers when appropriately rescaled. The ratio $\Omega/\lambda$ remains approximately constant across Re, indicating a form of universality in the core of the FTLE distribution.
Although we have studied the predictability problem from the point of view of the exponential growth of infinitesimal perturbations, it would be very interesting to investigate in detail the complementary regime of large errors and the statistics of finite-size Lyapunov exponents (Boffetta & Musacchio, 2017). Recent results have been obtained in the case of a decaying SPE cascade (Valade et al., 2024), where the authors were able to connect the hyperdiffusive behavior of Lagrangian fluid parcels with the anomalous diffusion of the system.
Acknowledgement
Supported by Italian Research Center on High Performance Computing Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2 within the activities of Spoke 3 (Astrophysics and Cosmos Observations). We acknowledge HPC CINECA for computing resources within the INFN-CINECA Grants INFN24-FieldTurb and INFN25-FieldTurb.